OpenStudy (anonymous):

integrate 2/2x^2+3x+1

OpenStudy (anonymous):

is it \[ \frac{2}{2x^2} + 3x + 1 = \frac{1}{x^2} + 3x + 1 \] ? i'm not sure how to read this.

OpenStudy (abb0t):

looks like you should be using partial fraction decomposition.

OpenStudy (abb0t):

Try factoring out the denominator so that PFD is easier.